OpenStudy (anonymous):

Anyone feeling generous enough to give a medal so I can lvl up? lol (gotta run to class in 3 mins >.<)

OpenStudy (anonymous):

<.<

OpenStudy (anonymous):

sure when u get out of class can you help me with a calculus problem??

OpenStudy (anonymous):

whats the problem? post it here :)

OpenStudy (anonymous):

if its straight forward, it shouldnt take too long.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

i dont want to go to American Lit :(

OpenStudy (anonymous):

problem 4 part b. your welcome! Thank u!! haha hated that class!

myininaya (myininaya):

dont go those classes are so boring

myininaya (myininaya):

i had this one class i kept not going to it and some how i got a B lol

OpenStudy (anonymous):

oh god, one of the branches of math im weak at :(

myininaya (myininaya):

which one?

OpenStudy (anonymous):

haha! problem 4 part b

OpenStudy (anonymous):

multivariable calculus.

OpenStudy (anonymous):

yea! its rocking my world!

myininaya (myininaya):

eww that looks bad lol

OpenStudy (anonymous):

i gotta run >.< if its not answered when my class is over, i'll work it out.

OpenStudy (anonymous):

thank u!!!!!! good luck in american lit! :)

myininaya (myininaya):

what is the non cal 1 way?

myininaya (myininaya):

i got e^t(2x-1) using \[\frac{df}{dt}=\frac{\delta f}{\delta x}\frac{dx}{dt}+\frac{\delta f}{\delta y}\frac{dy}{dt}\]

OpenStudy (anonymous):

i got 2x(e^t)+1(-e^-t) using the same formula. I must have done something wrong

OpenStudy (anonymous):

δf/δx=2x dx/dt=e^t δf/δy=1 dy/dt=-e^-t

myininaya (myininaya):

no you did right i just factored lol

OpenStudy (anonymous):

oh so I can leave it in 2x(e^t)+-e^-t form and it still be correct? and BTW I have no clue what the calc 1 one way is!

myininaya (myininaya):

lol that looks cal 1 to me i'm not sure what way they are looking for

OpenStudy (anonymous):

haha! then do you know what they are talking ab in part a?